# LIPM-Guided Reinforcement Learning for Stable and Perceptive Locomotion in Bipedal Robots

### v1

@@ IV-C Reward Design
$\begin{split}\boldsymbol{r}_{\text{stable}}&=\exp{(\|\hat{\boldsymbol{p}}_{\text{com}}-\boldsymbol{p}_{\text{com}}\|_{2}+(z_{c}-z)+\dot{\theta}_{\text{roll}}+\dot{\theta}_{\text{pitch}})}\\ &=\exp{(\boldsymbol{p}_{\text{e}}+z_{\text{e}}+\|\boldsymbol{\omega}_{\text{e}}\|_{1})}\end{split}$ (4)
where $\boldsymbol{p}_{e}$ denotes the CoM tracking error $\|\hat{\boldsymbol{p}}_{\text{com}}-\boldsymbol{p}_{\text{com}}\|_{2}$, $z_{e}$ represents the intercept error $(z_{c}-z)$, $\boldsymbol{\omega}_{e}=\left[\dot{\theta}_{roll},\dot{\theta}_{pitch}\right]^{T}$ is a vector include roll and pitch angular velocity.
Stable-Tracking Reward Fusion: For the perceptive locomotion challenge, when the bipedal robot loses stability during locomotion, it should slow down or stop to adjust its CoM position and torso orientation, which could be evaluated by (4), then resume velocity tracking. This highlights a fundamental priority: stability must take precedence over velocity tracking. However, traditional weighted-sum reward approaches fail to maintain this priority hierarchy effectively. Inspired by the RFM, a method that establishes task priorities through non-linear fusion of reward components, we apply this mechanism to combine stable rewards $r_{\text{stable}}\in(0,1]$ and velocity tracking rewards $r_{\text{vel}}\in(0,1]$:

### v2

@@ IV-C Reward Design
$\boldsymbol{r}_{\text{stable}}$ $=\exp\left(-\|\hat{\boldsymbol{p}}_{\text{com}}-\boldsymbol{p}_{\text{com}}\|_{2}-|z_{c}-z|\right$.
$\qquad\left.-(|\dot{\theta}_{\text{roll}}|+|\dot{\theta}_{\text{pitch}}|)\right)$
$=\exp\left(-\|\boldsymbol{p}_{\text{e}}\|_{2}-|z_{\text{e}}|-\|\boldsymbol{\omega}_{\text{e}}\|_{1}\right)$ (4)
where $\boldsymbol{p}_{e}=\hat{\boldsymbol{p}}_{\text{com}}-\boldsymbol{p}_{\text{com}}$ denotes the CoM tracking error vector, $z_{e}=z_{c}-z$ represents the intercept error, and $\boldsymbol{\omega}_{e}=\left[\dot{\theta}_{\text{roll}},\dot{\theta}_{\text{pitch}}\right]^{T}$ is the vector of roll and pitch angular velocities.
Stable-Tracking Reward Fusion: For the perceptive locomotion challenge, when the bipedal robot loses stability during locomotion, it should slow down or stop to adjust its CoM position and torso orientation, which could be evaluated by (IV-C), then resume velocity tracking. This highlights a fundamental priority: stability must take precedence over velocity tracking. However, traditional weighted-sum reward approaches fail to maintain this priority hierarchy effectively. Inspired by the RFM, a method that establishes task priorities through non-linear fusion of reward components, we apply this mechanism to combine stable rewards $r_{\text{stable}}\in(0,1]$ and velocity tracking rewards $r_{\text{vel}}\in(0,1]$: